The distance moved by a body along a straight line is proportional to the square root of the time . It's acceleration in terms of velocity V of the body is

Grade:11

1 Answers

Susmita
5 years ago
Given
x=t^\frac{1}{2}
Velocity v=dx/dt
                  =\frac{d}{dt} t^\frac{1}{2}=\frac{1}{2\sqrt{t}}
So acceleration
a=\frac{d}{dt}v=\frac{-1}{4t\sqrt{t}}
                     =\frac{-1}{2\sqrt{t}}(\frac{1}{2\sqrt{t}})^2×2
                     =-v.v2.2=-2v3
Hope this helps.If it does please approve the answer.